OpenStudy (anonymous):

Ethylene glycol (antifreeze) has a density of 1.11 g/cm3. What is the mass in grams of 387 mL of ethylene glycol?

OpenStudy (abmon98):

Density=Mass(g)/Volume(cm^3) 1.11=Mass(g)/387 ml 1.11*387=Mass(g)